A block of 3 kg is moving up on a ramp (by initial push). The angle of the ramp is 40 degrees. The coefficient of kinetic friction is 0.25. Calculate the acceleration of the object.
Answer #1

here,

mass of block , m = 3 kg

theta = 40 degree

the normal force , N = m * g * cos(theta)

uk = 0.25

the acceleratio , a = net force /effective mass

a = ( m *g * sin(theta) + uk * N /m


a = ( m *g * sin(theta) + uk * m * g * cos(theta)) /m

a = 9.81 * ( sin(40) + 0.25 * cos(40))

a = 8.18 m/s^2

the acceleration of system is 8.18 m/s^2 down the ramp